Session: Feb 2000 (session 1)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (morsiani at Sun Feb 27 12:22:56 MET 2000) Success

OAN\-Yebes (Paco Colomer at Mon Feb 28 10:14:41 MET 2000) Minor failures
No major problems and good weather. Tsys measurements wrong during the first hour of the experiment.

Effelsberg (A. Kraus at Mon Feb 28 15:21:21 MET 2000) Success

Onsala (A. Polatidis at Tue Feb 29 08:42:31 MET 2000) Success

Noto (C.Trigilio at Tue Mar 7 11:28:49 MET 2000) Failure
Error in the procedure file using drudg (option 21)-No recorder set for this esperiment. No data recorded. No tapes sent.

Calibration (Alastair Gunn at Wed Apr 19 13:08:11 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/el020c.antab for details